Can you please show me how to solve:
If a horse can run 47.5 miles in 60 minutes, how long will it take him to run .25 miles?
Thank you.....

Answer by Alan3354(69443)

If a horse can run 47.5 miles in 60 minutes, how long will it take him to run .25 miles?
--------------
The easiest way is a proportion.
47.5/60 = 0.25/t
47.5t = 60*0.25 = 15
t = 15/47.5 minutes
-----
Keep the units matching
miles/minutes = miles/minutes